find the derivative.
1. x = ln(x+y+1)
and
2. F(x) = (square root of quantity 1 + x)- ln(1 + square root of quantity x + 1)

Answer by stanbon(75887)

find the derivative.
1. x = ln(x+y+1)
1 = [1/(x+y+1)](1 + (y'))
y' = [x+y+1] -1
=============================
and
2. F(x) = sqrt(1 + x)- ln[1 + sqrt(x + 1)]
F'(x) = (1/2)(1+x)^(-1/2) - [1/(1+sqrt(x+1)] * [(1/2)(x+1)^(-1/2)]
====================================================================
Cheers,
Stan H.